OpenStudy (anonymous):

Find the sine, cosine, and tangent ratios of ZX. Help! I need someone to explain!!!😕☺️

OpenStudy (0_0youscareme):

i dont understand what you are asking can you please explain. do you mean the sine, cosine, and tan identities

OpenStudy (anonymous):

Umm yeah something like that. I only know how to solve using an angle. How do I solve this when it is asking me to use a line? ZX is a line right?

OpenStudy (0_0youscareme):

to make a line you need an equation for example y=x or y=2x. can you post the question word for word and ill see if i can help you

OpenStudy (anonymous):

Find the sine cosine and tangent ratios of ZX.

OpenStudy (0_0youscareme):

i think you are talking about the angle zxy. sin is opposite over hypotenuse. cos is adjacent over hypotenuse and tan is opposite over hypotenuse

OpenStudy (0_0youscareme):

i mean tan is opposite over adjacent

OpenStudy (0_0youscareme):

in this case 5 is the opposite side, 12 is the adjacent and 13 is the hypotenuse

OpenStudy (anonymous):

Yeah. I think so.

OpenStudy (0_0youscareme):

an easy way to remember the ratios is sohcahtoa. s stands for sin, o stands for opposite, h stands for hypotenuse, a stands for adjacent, c stands for cos, and t stands for tangent. think of it as (s=o/h)(c=a/h)(t=o/a)

OpenStudy (0_0youscareme):

0.o